# RanDumb: A Simple Approach that Questions the Efficacy of Continual Representation Learning

### v2

@@ Abstract
Continual learning has primarily focused on the issue of catastrophic forgetting and the associated stability-plasticity tradeoffs. However, little attention has been paid to the efficacy of continually learned representations, as representations are learned alongside classifiers throughout the learning process. Our primary contribution is empirically demonstrating that existing online continually trained deep networks produce inferior representations compared to a simple pre-defined random transforms. Our approach embeds raw pixels using a fixed random transform, approximating an RBF-Kernel initialized before any data is seen. We then train a simple linear classifier on top without storing any exemplars, processing one sample at a time in an online continual learning setting. This method, called RanDumb, significantly outperforms state-of-the-art continually learned representations across all standard online continual learning benchmarks. Our study reveals the significant limitations of representation learning, particularly in low-exemplar and online continual learning scenarios. Extending our investigation to popular exemplar-free scenarios with pretrained models, we find that training only a linear classifier on top of pretrained representations surpasses most continual fine-tuning and prompt-tuning strategies. Overall, our investigation challenges the prevailing assumptions about effective representation learning in online continual learning. Our code is available here.
@@ 1 Introduction
Figure 1: RanDumb projects raw pixels to a high dimensional space using random Fourier projections ($\varphi$), then decorrelate the features using Mahalanobis distance [42] and classify with the nearest class mean. The online update only involves updating a single sample covariance matrix and class-means.
@@
Building on the foundations of supervised deep learning, the prevalent approach in continual learning has been to jointly train representations alongside classifiers. This approach follows from the assumption that learned representations are expected to outperform fixed representation functions such as kernel classifiers, as demonstrated in supervised deep learning [34, 23, 57]. However, this assumption is never validated in continual learning, with scenarios having limited updates where networks might not be trained until convergence, such as online continual learning (OCL).
In this paper, we study the efficacy of representations derived from continual learning algorithms. Surprisingly, our findings suggest that these representations might not be as beneficial as presumed. To test this, we introduce a simple baseline method named RanDumb, which combines a random representation function with a straightforward linear classifier, illustrated in Figure 1. Our empirical evaluations, summarized in Table 1 (left, top), reveal that despite replacing the representation learning with a pre-defined random representation, RanDumb surpasses current state-of-the-art across latest online continual learning benchmarks [75].
We further expand our evaluations to scenarios that use pre-trained feature extractors [67]. By substituting our random projections with these feature extractors and retaining the linear classifier, RanDumb again outperforms leading methods as shown in Table 1 (right, top).
Design. RanDumb first projects input pixels into a high-dimensional space using a fixed kernel based on random Fourier basis, which is a low-rank data-independent approximation of the RBF Kernel [52]. Then, we use a simple linear classifier which first normalizes distances across different feature dimensions (anisotropy) with Mahalanobis distance [42] and then uses nearest class means for classification [44] (mechanism illustrated in Figure 2). In scenarios with pretrained feature extractors, we use the fixed pretrained model as embedder and learn a linear classifier as described above, similar to Hayes and Kanan [27].
@@ 2 RanDumb: Mechanism & Intuitions
$y_{\textrm{pred}}=\operatorname*{arg\,max}_{{i}\in\{1,\ldots,|C|\}}f({\bf x_{\textrm{test}}},\mathbf{\mu}_{i}),$ (1)
$\text{where}\quad f({\bf x_{\textrm{test}}},\mathbf{\mu}_{i}):={\bf x_{\textrm{test}}}^{\top}{\mathbf{\mu}_{i}}$ (2)
@@
Kernelization: Classes are typically not linearly separable in the pixel space, unlike in the feature space of deep models. Hence, we apply the kernel trick to embed the pixels in a better representation space, computing all distances between the data and class-means in this embedding space. This phenomena is illustrated on three toy examples to build intuitions in Figure 1 (right, Embed). We use an RBF-Kernel, which for two points $\mathbf{x}$ and $\mathbf{y}$ is defined as: $K_{\text{RBF}}(\mathbf{x},\mathbf{y})=\exp\left(-\gamma\|\mathbf{x}-\mathbf{y}\|^{2}\right)$ where $\gamma$ is a scaling parameter. However, calculating the RBF kernel is not possible due to the online continual learning constraints preventing computation of pairwise-distance between all points. Hence, we use a data-independent approximation, random Fourier projection $\phi(\mathbf{x})$, as given in [52]:
@@
We keep these $\omega$ bases fixed throughout online learning. Thus, our modified similarity function from Equation 1 is:
@@
Decorrelation: Projected raw pixels have feature dimensions with different variances (anisotropic). Hence, instead of naively computing ${\bf\phi(x_{\rm test})}^{\top}{\bar{\mathbf{\mu}}_{i}}$, we further decorrelate the feature dimensions using a Mahalonobis distance with the shrinked covariance matrix $\mathbf{S}$ using OAS shrinkage [15], inverse obtained by least squares minimization ($\mathbf{S}+\lambda\mathbf{I}$). We illustrate this phenomena as well on three toy examples in Figure 1 (right, Decorrelate) to build intuitions. Our similarity function finally is:
@@ 3 Experiments
We compare RanDumb with algorithms across online continual learning benchmarks with an emphasis on exemplar-free and low-exemplar storage regime.
Table 2: Overview of Benchmarks
@@
F (Wang et al. [67]) Many 1 All 0 No
Benchmarks. The benchmarks which we used in our experiments are summarized in Table 2. We aim for a comprehensive coverage and show results on four standard online continual learning benchmarks (A, B, D, E) which reflect the latest trends (’22-’24) across exemplar-free, contrastive-training [^2] Benchmark B is split into two sections: (B1) methods that do not rely on contrastive learning and heavy augmentation, and (B2) approaches that incorporate contrastive learning and extra augmentations., meta-continual learning, and network-expansion based approaches respectively. We also evaluate on a rehearsal-free offline continual learning benchmark C. These benchmarks are ordered by increasingly relaxed constraints, moving further away from the training scenario of RanDumb. Benchmark A closely matches RanDumb with one class per timestep and no stored exemplars. Benchmark B, D, E progressively relax the constraints on exemplars and classes per timestep. Benchmark C and E remove the online constraint by allowing unrestricted training and sample access within a task without exemplar-storage of past tasks.
We further test on exemplar-free scenarios in offline continual learning using Benchmark F [67] with the challenging one-class per task constraint borrowed from [75], i.e. testing over longer timespans. This benchmark allows using pretrained models along with unrestricted training time and access to all class samples at each timestep. However, RanDumb is restricted to learning from a single pass seeing only one sample at a time. RanDumb only learns a linear classifier over a given pretrained model in Benchmark F.
We use LAMDA-PILOT [61] codebase for all methods, except RanPAC and SLDA for which use their codebases. We use the original hyperparameters. We only change initial classes to 2 and number of classes per task to 1 and test using both ImageNet21K & ImageNet1K ViT-B/16 models.
@@
3.1 Results
Benchmark A. We assess continual learning models in the challenging setup of one class per timestep, closely mirroring our training assumptions, and present our results in Table 3. Comparing across rows, and see that RanDumb improves over prior state-of-the-art across all datasets with 2-6% margins. The only exception is PEC on CIFAR10, where RanDumb underperforms by 3.3%. Nonetheless, it outperforms the second-best model, GDumb with a 500 memory size, by 4.9%.
Benchmark B1. We present our results comparing with non-contrastive methods in Table 4. We notice that scenario allows two classes per task and relaxes the memory constraints for online continual learning methods, allowing for higher accuracies compared to Benchmark A. Despite that, RanDumb outperforms latest OCL algorithms on MNIST, CIFAR10 and CIFAR100—often by margins exceeding 10%. The lone exception is GDumb achieving a higher performance with 2K memory samples on TinyImageNet, indicating this already is in the high-memory regime.
Benchmark B2. We additionally compare our performance with the latest OCL approaches using contrastive losses with sophisticated data augmentations. As shown in in Table 5 (Left), these advancements provide large performance improvements over methods from Benchmark B.1. To compensate, we compare on lower exemplar budgets. The best approach, OnPro [68], outperforms RanDumb on CIFAR10 by 2.2% and TinyImageNet by 0.3%, but falls significantly short on CIFAR100 by 5.9%. Overall, RanDumb achieves strong results compared to representation learning using state-of-the-art contrastive learning approaches customized to continual learning, despite storing no exemplars.
Benchmark C. We compare against offline rehearsal-free continual learning approaches in Table 5 (Right) on CIFAR100. Despite online training, RanDumb outperforms PredKD by over 4% margins.
Benchmark D. We compare performance of RanDumb against meta-continual learning methods, which require large exemplars with buffer sizes of 1K in Table 6 (left). RanDumb achieves strong performance under these conditions, exceeding all prior work by a large margin of 9.1% on CIFAR100 and outperforms all but VR-MCL approach on the TinyImageNet dataset. GDumb performs the best on CIFAR10, indicating this is already in a large-exemplar regime uniquely unsuited for RanDumb.
Benchmark E. We compare RanDumb against network expansion-based online continual learning methods in Table 6 (right). These approaches grow model capacity to mitigate forgetting while dealing with shifts in the data distribution, and are allowed larger memory buffers. RanDumb matches the performance of the state-of-the-art method SEDEM [74] on MNIST, while exceeding it by 0.3% on CIFAR10 and 3.8% on CIFAR100.
@@ 3.1 Results
Table 8: Benchmark F We compare RanDumb with prompt-tuning approaches using ViT-B/16 ImageNet-21K/1K pretrained models using 2 init classes and 1 class per task setting. Most prompt-tuning based methods collapse, RanDumb achieves either state-of-the-art or second-best performance. RanPAC-imp is an improved version of the RanPAC mitigating the instability issues.
@@
3.2 Analysis of RanDumb
Comparisons across Architectures. In table 7 (right), we compare whether using random Fourier features as embeddings outperforms models across various architectures for continual representation learning. We use experience replay (ER) baseline in the task-incremental CIFAR100 setup (for details, see Mirzadeh et al. [45] as it differs significantly from earlier setups). Our comparison spanned various architectures. The findings revealed that RanDumb surpassed the performance of nearly all considered architectures, and achieved close to 94% of the joint multi-task performance. This suggests that RanDumb outperforms continual representation learning across architectures.
4 Related Work & Equivalent Formulations

### v3

# Random Representations Outperform Online Continually Learned Representations

@@ Abstract
Continual learning has primarily focused on the issue of catastrophic forgetting and the associated stability-plasticity tradeoffs. However, little attention has been paid to the efficacy of continually learned representations, as representations are learned alongside classifiers throughout the learning process. Our primary contribution is empirically demonstrating that existing online continually trained deep networks produce inferior representations compared to a simple pre-defined random transforms. Our approach projects raw pixels using a fixed random transform, approximating an RBF-Kernel initialized before any data is seen. We then train a simple linear classifier on top without storing any exemplars, processing one sample at a time in an online continual learning setting. This method, called RanDumb, significantly outperforms state-of-the-art continually learned representations across all standard online continual learning benchmarks. Our study reveals the significant limitations of representation learning, particularly in low-exemplar and online continual learning scenarios. Extending our investigation to popular exemplar-free scenarios with pretrained models, we find that training only a linear classifier on top of pretrained representations surpasses most continual fine-tuning and prompt-tuning strategies. Overall, our investigation challenges the prevailing assumptions about effective representation learning in online continual learning.
Figure 1: Our primary analysis in this work is ablating the deep feature extractor (bottom center) by replacing it with a random projection (top center) to isolate the effect of online continual representation learning in the deep network. We demonstrate that random projections not only match but consistently outperform continually learned representations, highlighting the poor quality of the continually learned representations. RanDumb (top) maps raw pixels to a high-dimensional space using random Fourier projections, then decorrelates the features using the Mahalanobis distance [43] and classifies based on the nearest class mean.
@@ 1 Introduction
Building on the foundations of supervised deep learning, the prevalent approach in continual learning has been to jointly train representations alongside classifiers. This approach simply follows from the assumption that learned representations are expected to outperform fixed representation functions such as kernel classifiers, as demonstrated in supervised deep learning [34, 23, 57]. However, this assumption is never validated in continual learning, with scenarios having limited updates where networks might not be trained until convergence, such as online continual learning (OCL).
In this paper, we study the efficacy of representations derived from continual learning algorithms. Surprisingly, our findings suggest that these representations might not be as beneficial as presumed. To test this, we introduce a simple baseline method named RanDumb, which combines a random representation function with a straightforward linear classifier, illustrated in detail in Figure 1. Our empirical evaluations, summarized in Table 1 (left, top), reveal that despite replacing the representation learning with a pre-defined random representation, RanDumb surpasses current state-of-the-art methods in latest online continual learning benchmarks [76].
We further expand our evaluations to scenarios incorporating methods that use pre-trained feature extractors [68]. By substituting our random projections with these feature extractors and retaining the linear classifier, RanDumb again outperforms leading methods as shown in Table 1 (right, top).
Design. RanDumb first projects input pixels into a high-dimensional space using a fixed kernel based on random Fourier basis, which is a low-rank data-independent approximation of the RBF Kernel [52]. Then, we use a simple linear classifier which first normalizes distances across different feature dimensions (anisotropy) with Mahalanobis distance [43] and then uses nearest class means for classification [44]. In scenarios with pretrained feature extractors, we use the fixed pretrained model as embedder and learn a linear classifier as described above, similar to Hayes and Kanan [27].
@@ 2 RanDumb: Mechanism & Intuitions
$y_{\textrm{pred}}=\operatorname*{arg\,max}_{{i}\in\{1,\ldots,|C|\}}f({\bf x_{\textrm{test}}},\mathbf{\mu}_{i}),\quad\text{where}\quad f({\bf x_{\textrm{test}}},\mathbf{\mu}_{i}):={\bf x_{\textrm{test}}}^{\top}{\mathbf{\mu}_{i}}$ (1)
@@
Kernelization: Classes are typically not linearly separable in the pixel space, unlike in the feature space of deep models. Hence, we randomly project the pixels into a high-dimensional representation space, computing all distances between the data and class-means in this embedding space. This phenomena is illustrated on three toy examples to build intuitions in Figure 2 (Embed). We use an RBF-Kernel, which for two points $\mathbf{x}$ and $\mathbf{y}$ is defined as: $K_{\text{RBF}}(\mathbf{x},\mathbf{y})=\exp\left(-\gamma\|\mathbf{x}-\mathbf{y}\|^{2}\right)$ where $\gamma$ is a scaling parameter. However, calculating the RBF kernel is not possible due to the online continual learning constraints preventing computation of pairwise-distance between all points. We use a data-independent approximation, random Fourier projection $\phi(\mathbf{x})$, as given in [52]:
@@
We keep these $\omega$ bases fixed throughout online learning. Thus, we obtain our modified similarity function from Equation 1 as:
@@
Decorrelation: Projected raw pixels have feature dimensions with different variances (anisotropic). Hence, instead of naively computing ${\bf\phi(x_{\rm test})}^{\top}{\bar{\mathbf{\mu}}_{i}}$, we further decorrelate the feature dimensions using a Mahalonobis distance with the shrinked covariance matrix $\mathbf{S}$ using OAS shrinkage [15], inverse obtained by least squares minimization ($\mathbf{S}+\lambda\mathbf{I}$). We illustrate this phenomena as well on three toy examples in Figure 2 (Decorrelate) to build intuitions. Our similarity function finally is:
@@ 3 Experiments
We compare RanDumb with algorithms across online continual learning benchmarks with an emphasis on exemplar-free and low-exemplar storage regime. All numbers in tables with the caption (Ref: table and citation) except our method are taken from the aforementioned table in the cited paper.
@@
F (Wang et al. [68], modified) Many 1 All 0 No
Benchmarks. The benchmarks which we used in our experiments are summized in Table on the right. We aim for a comprehensive coverage and show results on four standard online continual learning benchmarks (A, B, D, E) which reflect the latest trends (’22-’24) across exemplar-free, contrastive-training [^2] Benchmark B is split into two sections: (B1) methods that do not rely on contrastive learning and heavy augmentation, and (B2) approaches that incorporate contrastive learning and extra augmentations., meta-continual learning, and network-expansion based approaches respectively. We also evaluate on a rehearsal-free offline continual learning benchmark C. These benchmarks are ordered by increasingly relaxed constraints, moving further away from the training scenario of RanDumb.Benchmark A closely matches RanDumb with one class per timestep and no stored exemplars. Benchmark B, D, E progressively relax the constraints on exemplars and classes per timestep. Benchmark C and E remove the online constraint by allowing unrestricted training and sample access within a task without exemplar-storage of past tasks. Benchmark F allows using large pretrained models, modified by us with one class per task, i.e. testing learning over longer timespans.
@@
We further test on exemplar-free scenarios in offline continual learning using Benchmark F [68] with the challenging one-class per task constraint borrowed from [76]. This benchmark allows using pretrained models along with unrestricted training time and access to all class samples at each timestep. However, RanDumb is restricted to learning from a single pass seeing only one sample at a time. RanDumb only learns a linear classifier over a given pretrained model in Benchmark F.
We use LAMDA-PILOT [61] codebase for all methods, except RanPAC and SLDA for which use their codebases. We use the original hyperparameters. We only change initial classes to 2 and number of classes per task to 1 and test using both ImageNet21K and ImageNet1K ViT-B/16 models.
@@
3.1 Results
Benchmark A (single-class per task). We assess continual learning models in the challenging setup of one class per timestep, closely mirroring our training assumptions, and present our results in Table 3. Comparing across rows, and see that RanDumb improves over prior state-of-the-art across all datasets with 2-6% margins. The only exception is PEC on CIFAR10, where RanDumb underperforms by 3.3%. Nonetheless, it outperforms the second-best model, GDumb with a 500 memory size, by 4.9%.
@@ 3.1 Results
Benchmark B.1 (many-classes per task). We present our results comparing with non-contrastive methods in Table 3. We notice that scenario allows two classes per task and relaxes the memory constraints for online continual learning methods, allowing for higher accuracies compared to Benchmark A. Despite that, RanDumb outperforms latest OCL algorithms on MNIST, CIFAR10 and CIFAR100—often by margins exceeding 10%. The lone exception is GDumb achieving a higher performance with 2K memory samples on TinyImageNet, indicating that this already is in the high-memory regime.
Benchmark B.2 (many-classes per task, with contrastive losses and data augmentations). We additionally compare our performance with the latest OCL approaches using contrastive losses with sophisticated data augmentations. As shown in in Table 4 (Left), these advancements provide large performance improvements over methods from Benchmark B.1. To compensate, we compare on lower exemplar budgets. The best approach, OnPro [69], outperforms RanDumb on CIFAR10 by 2.2% and TinyImageNet by 0.3%, but falls significantly short on CIFAR100 by 5.9%. Overall, RanDumb achieves strong results compared to representation learning using state-of-the-art contrastive learning approaches customized to continual learning, despite storing no exemplars.
Benchmark C (rehearsal-free). We compare against offline rehearsal-free continual learning approaches in Table 4 (Right) on CIFAR100. Despite online training, RanDumb outperforms PredKD by over 4% margins.
Benchmark D (meta-continual learning). We compare performance of RanDumb against meta-continual learning methods, which require large exemplars with buffer sizes of 1K in Table 5 (left). RanDumb achieves strong performance under these conditions, exceeding all prior work by a large margin of 9.1% on CIFAR100 and outperforms all but VR-MCL approach on the TinyImageNet dataset. GDumb performs the best on CIFAR10, indicating this is already in a large-exemplar regime uniquely unsuited for RanDumb.
Benchmark E (network-expansion). We compare RanDumb against network expansion-based online continual learning methods in Table 5 (right). These approaches grow model capacity to mitigate forgetting while dealing with shifts in the data distribution, and are allowed larger memory buffers. RanDumb matches the performance of the state-of-the-art method SEDEM [75] on MNIST, while exceeding it by 0.3% on CIFAR10 and 3.8% on CIFAR100.
3.2 Analysis of RanDumb
@@
Comparisons across Architectures. In table 6 (right), we compare whether using random Fourier features as embeddings outperforms models across various architectures for continual representation learning. We use experience replay (ER) baseline in the task-incremental CIFAR100 setup
Figure 4: In previous experiments, models were trained from scratch, so we used random projection. Here, with a pretrained backbone, RanDumb starts with the frozen pretrained backbone to explore whether continual representation learning is necessary. By comparing this frozen backbone (RanDumb) with a continually trained one, we show that using the pretrained features consistently matches the best continually learned representations, similarly challenging the value of continual representation learning.
Table 7: Benchmark F We compare RanDumb with prompt-tuning approaches using ViT-B/16 ImageNet-21K/1K pretrained models using 2 init classes and 1 class per task setting. Most prompt-tuning based methods collapse and RanDumb achieves either state-of-the-art or second-best performance. RanPAC-imp is an improved version of the RanPAC mitigating the instability issues identified in a previous version of this work.
@@
(for details, see Mirzadeh et al. [45] as it differs significantly from earlier setups). Our comparison spanned various architectures. The findings revealed that RanDumb surpassed the performance of nearly all considered architectures, and achieved close to 94% of the joint multi-task performance. This suggests that RanDumb outperforms continual representation learning across architectures.
3.3 Should we learn representations when strong pre-trained features are available?
Say for a specific application (e.g., where the test data distribution is more or less known during training), practitioners should use strong pretrained models as a starting point as they are likely to perform better. However, we still ask the key question of whether representation learning is necessary by fixing the pretrained backbone and only training the linear classifier, as illustrated in Figure 4 in the next benchmark.
@@ 3.3 Should we learn representations when strong pre-trained features are available?
Why is the performance of prompting methods so low? We argue that the true test of continual learning lies in the ability to learn over prolonged periods. Our observations indicate that prompting methods collapse early across tasks because the learned prompts do not generalize effectively. Supporting this, parallel work [62] suggests that most current prompt-tuning methods lack prompt diversity and can be characterized by a single prompt, making classification performance heavily reliant on the quality of that prompt. We hypothesize when prompts are designed for a large number of classes (e.g., 20 or 50), they produce generally discriminative representations that extend to future tasks. However, prompts designed for only two classes, as in our case, have limited discriminative power, leading to the collapse of prompt-tuning methods across tasks.
@@
4 Related Works
@@ 5 Discussion and Concluding Remarks
Social Impact. RanDumb is an algorithm solely designed to perform a scientific study and we do not recommend use of RanDumb for any application in real-world production systems, hence no direct societal impact or explicit limitations on use in production systems is discussed.
@@ Acknowledgements
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper along with important assumptions.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: Limitation Section is provided in the supplementary material.
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: Rahimi and Recht [52] from our references details the theory for why random fourier representations perform so well quite beautifully. The random representations do not change (no continual aspect), hence the theory can be applied as-is in our case with no changes. We do not claim any novel theoretical contributions.
  13. 4.
    Experimental Result Reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: RanDumb is fairly simple to implement. We dedicated half a page towards explaining hyperparameters and other information needed to reproduce all of our results.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: We provide code in the supplementary material to reproduce our results.
  21. 6.
    Experimental Setting/Details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: We use standard datasets and splits, we provide hyperparameters in experimental details along with ablations in experiment sections to understand the contribution of each component in our algorithm.
  25. 7.
    Experiment Statistical Significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: We state here that the only random component being the random fourier features kernel across, otherwise our method is simple and exactly reproducible. We conducted experiments with three different initialisations corresponding to seeds of the random kernel in sklearn to investigate this and different kernel initialisations lead to around $\pm$0.2 variation in the reported results.
  29. 8.
    Experiments Compute Resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: Detailed in Section 3 in Implementation Details. For easy access, we restate: All experiments were conducted on a CPU server with a153 48-core Intel Xeon Platinum 8268 CPU and 392GB of RAM, requiring less than 30 minutes per experiment.
  33. 9.
    Code Of Ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: We have read the ethics guidelines and confirm that we do not use human subjects, use existing datasets, explicitly discuss social impacts.
  37. 10.
    Broader Impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: The primary contribution in RanDumb was not to introduce a novel state-of-the-art continual learning method, but challenge prevailing assumptions and open a discussion on the efficacy of representation learning in continual learning algorithms. As such, we do not recommend use of RanDumb for deployment in real-world production systems, hence no direct societal impact or explicit limitations on use in production systems is discussed.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: We do not have any high-risk model or dataset introduced.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: MNIST, CIFAR-10, CIFAR-100, tinyImageNet and miniImagenet are cited appropriately. The licenses for these datasets is not explicitly released, hence we do not include that information.
  49. 13.
    New Assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. Justification: Yes, we provide RanDumb with proper documentation under a GPL3 license.
  53. 14.
    Crowdsourcing and Research with Human Subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Justification: No crowdsourcing or research with human subjects was performed.
  57. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Justification: No crowdsourcing or research with human subjects was performed.
Appendix A Conceptual and Methodological Differences from GDumb [49]
Our core claim is that random representations from raw image pixels consistently outperform deep-learning-based representations designed for online continual learning. In contrast, GDumb’s central claim is that continual learning methods need not actually minimize forgetting of previous online samples as their performance can be entirely recovered simply by a baseline using the latest memory.
The key differences between thee two works is as follows:
  • Forgetting of Online Samples (GDumb): GDumb argues that continual learning methods suffer from forgetting online samples and addresses this by relying entirely on memory. This dismisses the value of directly learning from online data, as GDumb does not retain or utilize information from online samples. RanDumb, in contrast, exclusively learns from online samples without using memory, emphasizing the significance of ongoing data streams for performance.
  • Inadequate Representation Learning (RanDumb): GDumb’s approach to representation learning mirrors the experience replay (ER) baseline but does not address the quality of the learned representations. RanDumb explicitly focuses on the inadequacy of these representations and ablates their role to highlight their impact on continual learning. This reveals a key distinction in how each method evaluates representation quality in online learning settings.
Furthermore, the experimental setups for RanDumb and GDumb highlight their complementary nature:
  • Memory Settings: RanDumb primarily targets low-memory environments, whereas GDumb excels in high-memory scenarios. For example, in rehearsal-free settings without exemplar storage, a common trend in continual learning, GDumb would inherently produce random performance due to its dependence on memory. RanDumb, on the other hand, thrives in these low-memory contexts, providing an alternative solution when memory is constrained.
  • Complementary Nature: RanDumb and GDumb occupy complementary spaces within the continual learning landscape. RanDumb performs well in benchmarks where GDumb falters, and vice versa. As GDumb has been acknowledged as a valuable baseline, we argue that RanDumb similarly deserves recognition in the continual learning literature for its distinct strengths.
In summary, the only commonality between RanDumb and GDumb is that they serve as simple baselines. The points above underscore the fundamental distinctions between the two methods and the specific aims behind their development, as emphasized here and in the title of this work.
Appendix B Online Continual Learning: Our Setting
Current Problem formulation. We formally define the online continual learning (OCL) problem as follows. In classification settings, we aim to continually learn a function $f\colon\mathcal{X}\rightarrow\mathcal{Y}$, parameterized by $\theta_{t}$ at time $t$. OCL is an iterative process where each step consists of a learner receiving information and updating its model. For RanDumb, at each step $t$ of the interaction,
  1. 1.
    One data point $(x_{t},y_{t})\sim\pi_{t}$ sampled from a non-stationary distribution $\pi_{t}$ is revealed.
  2. 2.
    Learner updates the model $\theta_{t+1}$ using a compute budget, $B^{learn}_{t}$ and discards the datapoint.
Simplifications by Compared Approaches. Traditional online continual learning literature makes several concessions over this which makes the problem easier by allowing the datapoint to be saved for more timesteps. Training deep networks requires those simplifications as more data per batch helps stabilize the gradient updates. Typically, compared approaches store samples across for 10 timesteps, and performs an update with that batch of samples before discarding it. Most works further relax this by storing a memory buffer of samples indefinitely.
Drawbacks. Traditional online continual learning setups cannot effectively test for rapid adaptation because they use a class-incremental setup. Online learning is generally intended to enable quick adaptation to changing data and label distributions in a data stream. We believe a better formulation for online continual learning is described in [50].